# Course Allocation with Credits via Stable Matching

### v1

@@ 5.1 No lower quotas
We reduce from Min-MM, which is NP-complete, even if $G$ is a subdivision graph horton_1993_minimum . Let $G=(U,W,E)$ be a graph representing an instance of Min-MM, where $U$ and $W$ are the two sets of vertices, $|U|=n_{1}$ and $|W|=n_{2}$. Every vertex in $W$ has degree $2$. Let $K$ be the maximum size of a maximal matching. We build an instance $I$ of ca-ml, with students $S$ and courses $C$, as follows.
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, and $s_{i}^{3}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=2$, as well as $s_{i}^{4}$, with $T(s_{i}^{4})=1$. We also create courses $d_{i}^{1}$, $d_{i}^{2}$, and $d_{i}^{3}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$, and $O(d_{i}^{3})=1$. Students $s_{i}^{1}$, $s_{i}^{2}$, $s_{i}^{3}$ and $s_{i}^{4}$ belong to subset $S_{i}\subseteq S$, and courses $d_{i}^{1}$, $d_{i}^{2}$ and $d_{i}^{3}$ belong to subset $D_{i}\subseteq C$. For every $w_{i}\in W$, $s_{i}^{1}\in S^{1}\subseteq S$, $s_{i}^{2}\in S^{2}\subseteq S$, $s_{i}^{3}\in S^{3}\subseteq S$, and $s_{i}^{4}\in S^{4}\subseteq S$. Similarly, $d_{i}^{1}\in D^{1}\subseteq C$, $d_{i}^{2}\in D^{2}\subseteq C$, and $d_{i}^{3}\in D^{3}\subseteq C$.
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$, and a student $p_{j}$, with $T(p_{j})=1$. Additionally, we create a course $e$, $O(e)=2$. The upper quota of every course is $1$, except for $e$, which has upper quota $n_{2}-K$. For every $u_{j}\in U$, $c_{j}\in C^{\prime}\subseteq C$ and $p_{j}\in P\subseteq S$.
The master list of students (ML) is
$s_{1}^{1}\succ_{ML}s_{1}^{2}\succ_{ML}s_{2}^{1}\succ_{ML}s_{2}^{2}\succ_{ML}...\succ_{ML}s_{n_{2}}^{1}\succ_{ML}s_{n_{2}}^{2}\succ_{ML}\\ \succ_{ML}s_{1}^{3}\succ_{ML}s_{2}^{3}\succ_{ML}...\succ_{ML}s_{n_{2}}^{3}\succ_{ML}p_{1}\succ_{ML}p_{2}\succ_{ML}...\succ_{ML}p_{n_{1}}\succ_{ML}\\ \succ_{ML}s_{1}^{4}\succ_{ML}s_{2}^{4}\succ_{ML}...\succ_{ML}s_{n_{2}}^{4}$ (1)
@@
If in the Min-MM instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$, $s_{i}^{2}$ with $d_{i}^{1}$, $s_{i}^{3}$ with $d_{i}^{2}$ and $s_{i}^{4}$ with $d_{i}^{3}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, $s_{i}^{3}$ with $e$ and $s_{i}^{4}$ with $d_{i}^{3}$. Student $s_{i}^{3}$ contains courses $c_{j}$, $c_{k}$, and $d_{i}^{3}$ in her preference list to make sure that a course-complete first-coalition-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. For each student $p_{i}$, her preference list contains course $c_{j}$ only:
@@
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. Additionally, it includes student $p_{j}$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some agent from $S^{1}$, $S^{2}$ or $S^{3}$, otherwise it is matched with $p_{j}$.
@@
The preference lists of every $d_{i}^{1}\in D^{1}$, $d_{i}^{2}\in D^{2}$, and $d_{i}^{3}\in D^{3}$ are:
@@
Finally, the preference list of $e$ contains every student from $S^{3}$ in increasing subscript order. The students from $S^{3}$ matched to $e$ correspond to vertices in $G$ that are unmatched.
@@
$s_{1}^{3}$ $:d_{1}^{2}\succ_{s_{1}^{3}}d_{1}^{3}\succ_{s_{1}^{3}}c_{j}\succ_{s_{1}^{3}}c_{k}\succ_{s_{1}^{3}}e$
@@
Figure 6: Example of a Min-MM instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml
@@
$M^{\prime}$ is a first-coalition-stable matching.
@@
We prove the first-coalition stability of $M^{\prime}$ by showing that no student can belong to a first-blocking coalition. Every $s_{i}^{1}\in S^{1}$ is matched with either some $c_{j}\in C^{\prime}$, which is her most preferred course, or with $d_{i}^{1}$. As $O(d_{i}^{1})>O(c_{j})$, $(s_{i}^{1},\{c_{j}\})$ cannot form a first-blocking coalition. Similarly, every $s_{i}^{2}\in S^{2}$ is matched with either some $c_{k}\in C^{\prime}$, her most preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $c_{k}$ or with $d_{i}^{1}$, the same reasoning applies, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, so there is no blocking coalition containing $s_{i}^{2}$. Every $s_{i}^{4}\in S^{4}$ is matched with the only course in her preference list, $d_{i}^{3}$.
We now consider $s_{i}^{3}\in S^{3}$. This student is matched with either $d_{i}^{2}$ or with $e$. In the first case, she gets her most preferred course. In the second case, $(s_{i}^{2},d_{i}^{2})\in M$ and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched with some $w_{l}$ and $w_{m}$. Thus, there exist $s_{l}^{1}$ (or $s_{l}^{2}$) and $s_{m}^{1}$ (or $s_{m}^{2}$) such that $c_{j}$ and $c_{k}$ are matched with them. Thus, $s_{i}^{3}$ could form a first-blocking coalition with $d_{i}^{3}$ only. However, $O(d_{i}^{3})<O(e)$, so $(s_{i}^{3},\{d_{i}^{3}\})$ is not a first-blocking coalition. Finally, every $p_{j}\in P$ is either matched with $c_{j}$, the only course in her preference list, or she has no partner. In this second case, $c_{j}$ is matched with some $s_{i}^{1}$ or $s_{i}^{2}$, so $c_{j}$ prefers its current student over $p_{j}$, and thus $p_{j}$ cannot be part of a first-blocking coalition. ∎
Now, assume that there exists a course-complete first-coalition-stable matching $M^{\prime}$ in $I$. We now show how to obtain a maximal matching $M$ of size at most $K$ in $G$. First, as $M^{\prime}$ is course-complete, $|S_{M^{\prime}}(e)|=q^{+}(e)=n_{2}-K$, so there are $n_{2}-K$ students $s_{i}^{3}\in S^{3}$ matched to $e$. For each pair $(s_{i}^{3},e)\in M^{\prime}$, $s_{i}^{3}$ would prefer to be matched with $d_{i}^{2}$. Thus, because $M^{\prime}$ is first-coalition-stable, we must have that $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$, which in turn implies that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$. Likewise, $s_{i}^{3}$ prefers to be matched with $d_{i}^{3}$ and with either $c_{j}$ or with $c_{k}$. Hence, because $M^{\prime}$ is first-coalition-stable, and since $d_{i}^{3}$ ranks $s_{i}^{3}$ first, both $c_{j}$ and $c_{k}$ must have better students than $s_{i}^{3}$, so students from $S^{1}$ or $S^{2}$ (or possibly $S^{3}$). From this we get that, due to $M^{\prime}$ being course-complete, $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. Thus, if $(s_{i}^{3},e)\in M^{\prime}$, no student from $S_{i}$ is matched with a course from $C^{\prime}$. These $S_{i}$ subsets correspond to $n_{2}-K$ vertices $w_{i}\in G$ that are not matched in the maximal matching $M$.
Let us now consider the $K$ remaining $S_{i}$ subsets. For some of these subsets, there is no $s_{i}^{l}\in S_{i}$ and $c_{j}\in C^{\prime}$ such that $(s_{i}^{l},c_{j})\in M^{\prime}$. In these cases, we have that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{3},d_{i}^{3})\in M^{\prime}$, as $M^{\prime}$ is first-coalition-stable and course-complete, so every $d_{i}^{l}\in D_{i}$ must have one student. In these cases, we have that both $c_{j}$ and $c_{k}$ are matched with students from $S^{1}$, $S^{2}$ or $S^{3}$, to prevent $s_{i}^{3}$ forming a first-blocking coalition with either $c_{j}$ or $c_{k}$. For the remaining subsets $S_{i}$, there exists exactly one pair of the form $(s_{i}^{l},c_{j})$ in $M^{\prime}$.
@@
If $(s_{i}^{l},c_{j})\in M^{\prime}$ for some $s_{i}^{l}\in S_{i}$, then $M^{\prime}$ includes no other pair between a member of $S_{i}$ and a member of $C^{\prime}$.
@@
If $s_{i}^{1}$ is matched with $c_{j}$, then $(s_{i}^{2},d_{i}^{1})$ must be in $M^{\prime}$, otherwise $d_{i}^{1}$ has no students and is not course-complete. Thus, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. If $s_{i}^{2}$ is matched with $c_{k}$, then, by the same reasoning, $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. Finally, if $(s_{i}^{3},c_{j})\in M^{\prime}$ (or $(s_{i}^{3},c_{k})\in M^{\prime}$), then, to preserve first-coalition stability, $(s_{i}^{3},d_{i}^{3})\in M^{\prime}$, $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, and $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$. ∎
Hence, for every $S_{i}\subseteq S$, either one $s_{i}^{l}\in S_{i}$ is matched with some $c_{j}$ or none are. If some $s_{i}^{l}\in S_{i}$ is matched with some $c_{j}$, add $(u_{j},w_{i})$ to the matching $M$ in $G$, which is a valid matching because every vertex is matched with at most one, adjacent, vertex. As there are at least $n_{2}-K$ subsets where no member is matched with some $c_{j}$, we have that there are at most $K$ $S_{i}$ subsets where exactly one member is matched with one $c_{j}$. This means that the matching $M$ is of size at most $K$. We now show that $M$ is maximal.
@@
Looking for a contradiction, assume that it is not, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other. This implies that in $M^{\prime}$, $s_{i}^{3}$ is matched with either $e$ or $d_{i}^{3}$. In the first case, $c_{j}$ is matched with $p_{j}$, and $d_{i}^{3}$ is matched with $s_{i}^{4}$. However, $s_{i}^{3}$ prefers $\{d_{i}^{3},c_{j}\}$ over $e$, $c_{j}$ prefers $s_{i}^{3}$ over $p_{j}$, and $d_{i}^{3}$ prefers $s_{i}^{3}$ over $s_{i}^{4}$. Thus, $(s_{i}^{3},\{d_{i}^{3},c_{j}\})$ is a first-blocking coalition.
In the second case, $c_{j}$ is matched with $p_{j}$. However, $s_{i}^{3}$ prefers $c_{j}$ over being unmatched, as $O(d_{i}^{3})=1$ and $T(s_{i}^{3})=2$, and $c_{j}$ prefers $s_{i}^{3}$ over $p_{j}$. Thus, $(s_{i}^{3},\{c_{j}\})$ is a first-blocking coalition. In either case we get a contradiction to the stability of $M^{\prime}$. Therefore, $M$ is maximal. ∎
Hence, there exists a maximal matching of size at most $K$ in $G$ if and only if there exists a course-complete first-coalition-stable matching in $I$. As Min-MM is NP-complete, so is determining the existence of a course-complete first-coalition-stable matching in a ca-ml instance. ∎
@@
Determining whether a ca-ml instance has a course-complete coalition-stable matching is NP-hard, even if every student takes up to $2$ credits, every course has $1$ or $2$ credits, and each student’s preference list is of length at most $5$.
@@
As an aside, one can determine whether a ca-ml instance $I$ has a student-complete first-coalition-stable matching in polynomial time, as we now show.
Proposition 3.
Determining whether a ca-ml instance $I$ has a student-complete first-coalition-stable (feasible) matching is solvable in polynomial time.
@@
In order to find a student-complete first-coalition-stable matching, or show that none exists, we use a variant of Serial Dictatorship. In this variant, starting from the beginning of the preference list, for each student $s_{i}$ we attempt to find a subset $C_{i}\subseteq P(s_{i})$ such that $s_{i}$ can be matched with $C_{i}$ (i.e. the resulting matching is first-coalition stable and no constraints are violated) and $O(C_{i})=T(s_{i})$. If such a $C_{i}$ exists, then it is unique. Assume the opposite, so that $s_{i}$ can be matched with courses from subsets $C_{i}\subseteq P(s_{i})$ and $C_{i}^{\prime}\subseteq P(s_{i})$, such that $C_{i}\neq D_{i}$, $O(C_{i})=O(D_{i})=T(s_{i})$, both $C_{i}$ and $D_{i}$ are feasible, and every course in both $C_{i}$ and $D_{i}$ has space for $s_{i}$. Let $c_{i}$ be $s_{i}$’s most preferred course in $C_{i}\backslash D_{i}$ and $d_{i}$ her most preferred course in $D_{i}\backslash C_{i}$. Without loss of generality, assume that $c_{i}\succ_{s_{i}}d_{i}$. Then $s_{i}$ prefers $C_{i}$ over $D_{i}$, and so if matched with the courses in $D_{i}$ she will form a first-blocking coalition with a subset of courses of $C_{i}$. Hence there is at most one possible $C_{i}$ that satisfies all conditions.
The algorithm proceeds as follows. First, the first student in the master list, $s_{1}$, selects the unique subset $C_{1}$, should it exist, and they are matched together. Then $s_{2}$ selects the unique subset $C_{2}$, and so on. If for some $s_{i}$ such a $C_{i}$ does not exist, then no student-complete first-coalition-stable matching exists, because $s_{i}$ cannot choose a subset of courses that allows her to take her maximum number of credits, and the previous students in the master list cannot choose different courses, which may have freed some space in some $c_{j}\in P(s_{i})$ and allowed $s_{i}$ to select a different subset. Otherwise, for every $s_{i}$, $O(C_{i})=T(s_{i})$, so we have found a student-complete first-coalition-stable matching. ∎
We now show that determining the existence of a student-complete (or course-complete) pair-size-stable matching is NP-hard. In order to prove this result, we will reduce from Exact-MM. In this problem, we have a graph $G$ and a positive integer $K$, and the goal is to determine whether there exists a maximal matching of size exactly $K$. This problem is NP-complete, even if $G$ is a subdivision graph iwama_2002_stable ; manlove_2002_hard (manlove_algorithmics_2013, , page 133), which implies that it is bipartite and all vertices in one of the sides have degree $2$.
Theorem 8.
It is NP-hard to determine whether an instance of ca-ml has a student-complete (or course-complete) pair-size-stable matching, even if both students and courses are ordered according to master lists, every student takes up to $2$ credits, and every course has $1$ or $2$ credits.
Proof.
@@
We reduce from Exact-MM, which is NP-complete, even if $G$ is bipartite and all vertices in one of the sides have degree $2$. Let $G=(U,W,E)$ be a graph representing an instance of Exact-MM, where $U$ and $W$ are the two sets of vertices, $|U|=n_{1}$ and $|W|=n_{2}$. Every vertex in $W$ has degree $2$. Our goal is to find a maximal matching of size $K$. We build an instance $I$ of ca-ml, with students $S$ and courses $C$, as follows.

### v2

@@ 5.1 No lower quotas
We reduce from min-mm, which is NP-complete, even if $G$ is a subdivision graph Horton and Kilakos (1993). Let $G=(U,W,E)$ be a graph and let $K$ be an integer representing an instance of min-mm, where $U$ and $W$ are the two sets of vertices in the bipartition, $|U|=n_{1}$ and $|W|=n_{2}$, and $K$ is the target size of a maximal matching. Without loss of generality assume that every vertex in $W$ has degree $2$. We build an instance $I$ of ca-ml, with students $S$ and courses $C$, as follows.
For every $w_{i}\in W$, we create the following students: $s_{i}^{1}$, $s_{i}^{2}$, and $s_{i}^{3}$, with $T(s_{i}^{1})=T(s_{i}^{2})=T(s_{i}^{3})=2$, as well as $s_{i}^{4}$, with $T(s_{i}^{4})=1$. We also create courses $d_{i}^{1}$, $d_{i}^{2}$, and $d_{i}^{3}$, with $O(d_{i}^{1})=O(d_{i}^{2})=2$, and $O(d_{i}^{3})=1$. Create the following subsets: $S_{i}=\{s_{i}^{1},s_{i}^{2},s_{i}^{3},s_{i}^{4}\}$ ($1\leq i\leq n_{2}$), $D_{i}=\{d_{i}^{r}:1\leq i\leq n_{2}\wedge 1\leq r\leq 3\}$, $S^{r}=\{s_{i}^{r}:1\leq i\leq n_{2}\}$ ($1\leq r\leq 4$), and $D^{r}=\{d_{i}^{r}:1\leq i\leq n_{2}\}$ ($1\leq r\leq 3$).
For every $u_{j}\in U$, we create a course $c_{j}$, with $O(c_{j})=1$, and a student $p_{j}$, with $T(p_{j})=1$. Additionally, we create a course $e$, $O(e)=2$. The upper quota of every course is $1$, except for $e$, which has upper quota $n_{2}-K$. Let $C^{\prime}=\{c_{j}:1\leq j\leq n_{1}\}$ and $P=\{p_{j}:1\leq j\leq n_{1}\}$.
The master list of students $(\succ_{MLS})$ is
$s_{1}^{1}\succ_{MLS}s_{1}^{2}\succ_{MLS}s_{2}^{1}\succ_{MLS}s_{2}^{2}\succ_{MLS}\dots\succ_{MLS}s_{n_{2}}^{1}\succ_{MLS}s_{n_{2}}^{2}\succ_{MLS}\\ s_{1}^{3}\succ_{MLS}s_{2}^{3}\succ_{MLS}\dots\succ_{MLS}s_{n_{2}}^{3}\succ_{MLS}p_{1}\succ_{MLS}p_{2}\succ_{MLS}\dots\succ_{MLS}p_{n_{1}}\succ_{MLS}\\ s_{1}^{4}\succ_{MLS}s_{2}^{4}\succ_{MLS}\dots\succ_{MLS}s_{n_{2}}^{4}$
@@
If in the min-mm instance $w_{i}$ is matched with $u_{j}$, then $s_{i}^{1}$ is to be matched with $c_{j}$, $s_{i}^{2}$ with $d_{i}^{1}$, $s_{i}^{3}$ with $d_{i}^{2}$ and $s_{i}^{4}$ with $d_{i}^{3}$. Similarly if $w_{i}$ is matched with $u_{k}$, then $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $c_{k}$, $s_{i}^{3}$ with $d_{i}^{2}$ and $s_{i}^{4}$ with $d_{i}^{3}$. If, however, $w_{i}$ is unmatched, $s_{i}^{1}$ is to be matched with $d_{i}^{1}$, $s_{i}^{2}$ with $d_{i}^{2}$, $s_{i}^{3}$ with $e$ and $s_{i}^{4}$ with $d_{i}^{3}$. Student $s_{i}^{3}$ contains courses $c_{j}$, $c_{k}$, and $d_{i}^{3}$ in her preference list to make sure that a course-complete first-coalition-stable matching in $I$ corresponds to a maximal matching in $G$, so that if $s_{i}^{3}$ is matched with $e$, some students from $S^{1}$ or $S^{2}$ are matched with $c_{j}$ and $c_{k}$. For each student $p_{j}\in P$, her preference list contains course $c_{j}$ only:
@@
The preference list of every $c_{j}\in C^{\prime}$ depends on the vertices adjacent to $u_{j}$ in $G$. For every $w_{i}$ adjacent to $u_{j}$, $c_{j}$ includes in its preference list either $s_{i}^{1}$ or $s_{i}^{2}$, as well as $s_{i}^{3}$. Additionally, it includes student $p_{j}$. The order of all students is derived from the master list. If, for example, $u_{j}$ is adjacent to $w_{3}$, $w_{4}$ and $w_{5}$, the preference list of $c_{j}$ is given below. Here the superscript $l_{i}$ indicates a number that can be either $1$ or $2$. If $u_{j}$ is matched in $G$ then $c_{j}$ is matched with some student from $S^{1}$, $S^{2}$ or $S^{3}$, otherwise it is matched with $p_{j}$.
@@
The preference lists of every course $d_{i}^{1}\in D^{1}$, $d_{i}^{2}\in D^{2}$, and $d_{i}^{3}\in D^{3}$ are:
@@
Finally, the preference list of course $e$ contains every student from $S^{3}$ in increasing subscript order:
@@
The students from $S^{3}$ matched to $e$ correspond to vertices in $G$ that are unmatched.
@@
$s_{1}^{3}$ $:d_{1}^{2}\succ_{s_{1}^{3}}d_{1}^{3}\succ_{s_{1}^{3}}c_{1}\succ_{s_{1}^{3}}c_{2}\succ_{s_{1}^{3}}e$
@@
Figure 6: Example of a min-$m$ instance, with $K=2$, and some of the preference lists in the transformed instance of ca-ml
@@
$M^{\prime}$ is a first-coalition-stable matching in $I$.
@@
We prove the first-coalition stability of $M^{\prime}$ by showing that no student can belong to a first-blocking coalition. Every $s_{i}^{1}\in S^{1}$ is matched with either some $c_{j}\in C^{\prime}$, which is her most-preferred course, or with $d_{i}^{1}$. As $O(d_{i}^{1})>O(c_{j})$, $(s_{i}^{1},\{c_{j}\})$ cannot form a first-blocking coalition. Similarly, every $s_{i}^{2}\in S^{2}$ is matched with either some $c_{k}\in C^{\prime}$, her most-preferred course, with $d_{i}^{1}$, or with $d_{i}^{2}$. If $s_{i}^{2}$ is matched with $c_{k}$ or with $d_{i}^{1}$, the same reasoning applies as for $s_{i}^{1}$, and if she is matched with $d_{i}^{2}$, then $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, and $s_{i}^{1}\succ_{d_{i}^{1}}s_{i}^{2}$, so there is no blocking coalition containing $s_{i}^{2}$. Every $s_{i}^{4}\in S^{4}$ is matched with the only course in her preference list, $d_{i}^{3}$.
We now consider $s_{i}^{3}\in S^{3}$. This student is matched with either $d_{i}^{2}$ or with $e$. In the first case, she gets her most-preferred course. In the second case, $(s_{i}^{2},d_{i}^{2})\in M$, $s_{i}^{2}\succ_{d_{i}^{2}}s_{i}^{3}$, and, because $M$ is maximal in $G$, $u_{j}$ and $u_{k}$ are matched in $M$ with some $w_{p}$ and $w_{q}$ respectively. Thus, there exist $s_{p}^{l_{p}},s_{q}^{l_{q}}\in S^{1}\cup S^{2}$ such that $(s_{p}^{l_{p}},c_{j})\in M^{\prime}$ and $(s_{q}^{l_{q}},c_{k})\in M^{\prime}$. Thus, $s_{i}^{3}$ could form a first-blocking coalition with $d_{i}^{3}$ only. However, $O(d_{i}^{3})<O(e)$, so $(s_{i}^{3},\{d_{i}^{3}\})$ is not a first-blocking coalition. Finally, every $p_{j}\in P$ is either matched with $c_{j}$, the only course in her preference list, or she has no partner. In this second case, $c_{j}$ is matched with some $s_{i}^{1}$ or $s_{i}^{2}$, so $c_{j}$ prefers its current student over $p_{j}$, and thus $p_{j}$ cannot be part of a first-blocking coalition. ∎
Now, assume that there exists a course-complete first-coalition-stable matching $M^{\prime}$ in $I$. We now show how to obtain a maximal matching $M$ of size at most $K$ in $G$. We begin by establishing two lemmas that give useful properties of $M^{\prime}$.
@@
There are $n_{2}-K$ subsets $S_{i}$ for which $M^{\prime}$ includes no pair between a member of $S_{i}$ and a member of $C^{\prime}$.
@@
As $M^{\prime}$ is course-complete, $|S_{M^{\prime}}(e)|=q^{+}(e)=n_{2}-K$, so there are $n_{2}-K$ students $s_{i}^{3}\in S^{3}$ matched to $e$. For each pair $(s_{i}^{3},e)\in M^{\prime}$, $s_{i}^{3}$ would prefer to be matched with $d_{i}^{2}$. Thus, because $M^{\prime}$ is first-coalition-stable, we must have that $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$, which in turn implies that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$. Likewise, $s_{i}^{3}$ prefers to be matched with $d_{i}^{3}$ and with either $c_{j}$ or with $c_{k}$. Hence, because $M^{\prime}$ is first-coalition-stable, and since $d_{i}^{3}$ ranks $s_{i}^{3}$ first, both $c_{j}$ and $c_{k}$ must have better students than $s_{i}^{3}$, so students from $S^{1}$ or $S^{2}$ (or possibly $S^{3}$). We also infer that, due to $M^{\prime}$ being course-complete, $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. Thus, if $(s_{i}^{3},e)\in M^{\prime}$, no student from $S_{i}$ is matched with a course from $C^{\prime}$. These $S_{i}$ subsets correspond to $n_{2}-K$ vertices $w_{i}\in G$ that are not matched in the maximal matching $M$. ∎
Let us now consider the $K$ remaining $S_{i}$ subsets. For some of these subsets, there is no $s_{i}^{l_{i}}\in S_{i}$ and $c_{j}\in C^{\prime}$ such that $(s_{i}^{l_{i}},c_{j})\in M^{\prime}$. In these cases, we have that $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{3},d_{i}^{3})\in M^{\prime}$, as $M^{\prime}$ is first-coalition-stable and course-complete, so every $d_{i}^{l}\in D_{i}$ must have one student. In these cases, we have that both $c_{j}$ and $c_{k}$ are matched with students from $S^{1}$, $S^{2}$ or $S^{3}$, to prevent $s_{i}^{3}$ forming a first-blocking coalition with either $c_{j}$ or $c_{k}$. For the remaining subsets $S_{i}$, there exists exactly one pair of the form $(s_{i}^{l},c_{j})$ in $M^{\prime}$.
Lemma 5.
If $(s_{i}^{l_{i}},c_{j})\in M^{\prime}$ for some $s_{i}^{l_{i}}\in S_{i}$, then $M^{\prime}$ includes no other pair between a member of $S_{i}$ and a member of $C^{\prime}$.
Proof.
If $s_{i}^{1}$ is matched with $c_{j}$, then $(s_{i}^{2},d_{i}^{1})$ must be in $M^{\prime}$, otherwise $d_{i}^{1}$ has no students and $M^{\prime}$ is not course-complete. Thus, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. If $s_{i}^{2}$ is matched with $c_{k}$, then, by the same reasoning, $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, $(s_{i}^{3},d_{i}^{2})\in M^{\prime}$ and $(s_{i}^{4},d_{i}^{3})\in M^{\prime}$. Finally, if $(s_{i}^{3},c_{j})\in M^{\prime}$ (or $(s_{i}^{3},c_{k})\in M^{\prime}$), then, to preserve first-coalition stability, $(s_{i}^{3},d_{i}^{3})\in M^{\prime}$, $(s_{i}^{1},d_{i}^{1})\in M^{\prime}$, and $(s_{i}^{2},d_{i}^{2})\in M^{\prime}$. ∎
Hence, for every $S_{i}$ $(1\leq i\leq n_{2})$, Lemmas 4 and 5 imply that either one $s_{i}^{l_{i}}\in S_{i}$ is matched in $M^{\prime}$ with some $c_{j}\in C^{\prime}$ or no member of $S_{i}$ is matched in $M^{\prime}$ with a member of $C^{\prime}$. If some $s_{i}^{l_{i}}\in S_{i}$ is matched with some $c_{j}$ in $M^{\prime}$, add $(u_{j},w_{i})$ to $M$ in $G$, which is a valid matching in $I$ because every vertex is matched with at most one, adjacent, vertex. As there are at least $n_{2}-K$ subsets where no member is matched with some $c_{j}$, we have that there are at most $K$ $S_{i}$ subsets where exactly one member is matched with one $c_{j}$. This means that the matching $M$ is of size at most $K$. We now show that $M$ is maximal.
@@
Looking for a contradiction, assume that $M$ is not maximal, so that there exists unmatched $u_{j},w_{i}$ that are adjacent to each other in $G$. This implies that in $M^{\prime}$, $c_{j}$ is matched with $p_{j}$, and $s_{i}^{3}$ is matched with either $e$ or $d_{i}^{3}$. In the first case, $d_{i}^{3}$ is matched with $s_{i}^{4}$ in $M^{\prime}$. However, $s_{i}^{3}$ prefers $\{d_{i}^{3},c_{j}\}$ over $e$, $c_{j}$ prefers $s_{i}^{3}$ over $p_{j}$, and $d_{i}^{3}$ prefers $s_{i}^{3}$ over $s_{i}^{4}$. Thus, $(s_{i}^{3},\{d_{i}^{3},c_{j}\})$ is a first-blocking coalition of $M^{\prime}$.
In the second case, $s_{i}^{3}$ still has capacity for $c_{j}$, as $O(c_{j})=d_{i}^{3})=1$ and $T(s_{i}^{3})=2$, and $c_{j}$ prefers $s_{i}^{3}$ over $p_{j}$. Thus, $(s_{i}^{3},\{c_{j}\})$ is a first-blocking coalition of $M^{\prime}$. In either case we get a contradiction to the first-coalition stability of $M^{\prime}$ in $I$. Therefore, $M$ is maximal in $G$. ∎
Hence, there exists a maximal matching of size at most $K$ in $G$ if and only if there exists a course-complete first-coalition-stable matching in $I$. As min-mm is NP-complete, determining the existence of a course-complete first-coalition-stable matching in a ca-ml instance is NP-hard. ∎
@@
Determining whether a ca-ml instance has a course-complete coalition-stable matching is NP-hard, even if every student takes up to two credits, every course has one or two credits, and each student’s preference list is of length at most $5$.
@@
We now show that determining the existence of a student-complete (or course-complete) pair-size-stable matching is NP-hard. In order to prove this result, we will reduce from exact-mm. In this problem, we have a graph $G$ and a positive integer $K$, and the goal is to determine whether there exists a maximal matching of size exactly $K$. This problem is NP-complete, even if $G$ is a subdivision graph Manlove et al. (2002)(Manlove, 2013, page 133), which implies that it is bipartite and all vertices in one of the sides have degree $2$.
Theorem 8.
It is NP-hard to determine whether an instance of ca-ml has a student-complete (or course-complete) pair-size-stable matching, even if both students and courses are ordered according to master lists, every student takes up to two credits, and every course has one or two credits.
@@
We reduce from exact-mm, which is NP-complete, even if $G$ is bipartite and all vertices in one of the sides have degree $2$ Manlove et al. (2002)(Manlove, 2013, page 133). Let $G=(U,W,E)$ be a graph and $K$ be an integer representing an instance of exact-mm, where $U$ and $W$ are the two sets of vertices in the bipartition, $|U|=n_{1}$ and $|W|=n_{2}$, and $K$ is the target size of a maximal matching. Without loss of generality assume that every vertex in $W$ has degree $2$. We build an instance $I$ of ca-ml, with students $S$ and courses $C$, as follows.